\begin{thebibliography}{10}

\bibitem{DBLP:conf/sfm/KwiatkowskaNP07}
Kwiatkowska, M., Norman, G., Parker, D.:
\newblock Stochastic model checking.
\newblock  Proc.\ of SFM. Vol. 4486 of LNCS, Springer (2007)  220--270

\bibitem{DBLP:conf/lics/Kwiatkowska03}
Kwiatkowska, M.:
\newblock Model checking for probability and time: from theory to practice.
\newblock  {LICS}, IEEE CS (2003)  351

\bibitem{DBLP:series/natosec/Baier16}
Baier, C.:
\newblock Probabilistic model checking.
\newblock  Dependable Software Systems Engineering. Vol.~45 of {NATO} Science
  for Peace and Security Series - {D:} Information and Communication Security.
\newblock {IOS} Press (2016)  1--23

\bibitem{Kat16}
Katoen, J.P.:
\newblock The probabilistic model checking landscape.
\newblock  Proc. of LICS, ACM (2016)  31--46

\bibitem{DBLP:conf/atva/DehnertJWAK14}
Dehnert, C., Jansen, N., Wimmer, R., {\'{A}}brah{\'{a}}m, E., Katoen, J.P.:
\newblock Fast debugging of {PRISM} models.
\newblock  Proc.\ of ATVA. Vol. 8837 of LNCS, Springer (2014)  146--162

\bibitem{junges-et-al-tacas-2016}
Junges, S., Jansen, N., Dehnert, C., Topcu, U., Katoen, J.P.:
\newblock Safety-constrained reinforcement learning for {MDP}s.
\newblock  Proc.\ of TACAS. (2016)

\bibitem{dehnert-et-al-cav-2015}
Dehnert, C., Junges, S., Jansen, N., Corzilius, F., Volk, M., Bruintjes, H.,
  Katoen, J.P., {\'{A}}brah{\'{a}}m, E.:
\newblock Prophesy: {A} probabilistic parameter synthesis tool.
\newblock  Proc.\ of CAV. Vol. 9206 of LNCS, Springer (2015)  214--231

\bibitem{DBLP:conf/atva/QuatmannD0JK16}
Quatmann, T., Dehnert, C., Jansen, N., Junges, S., Katoen, J.P.:
\newblock Parameter synthesis for {M}arkov models: Faster than ever.
\newblock  Proc.\ of ATVA. Vol. 9938 of LNCS (2016)  50--67

\bibitem{DBLP:conf/safecomp/0001JK16}
Volk, M., Junges, S., Katoen, J.P.:
\newblock Advancing dynamic fault tree analysis - get succinct state spaces
  fast and synthesise failure rates.
\newblock  Proc.\ of SAFECOMP. Vol. 9922 of LNCS, Springer (2016)  253--265

\bibitem{DBLP:conf/atva/0001DKKW16}
Jansen, N., Dehnert, C., Kaminski, B.L., Katoen, J.P., Westhofen, L.:
\newblock Bounded model checking for probabilistic programs.
\newblock  Proc.\ of ATVA. Vol. 9938 of LNCS (2016)  68--85

\bibitem{mrmc}
Katoen, J.P., Zapreev, I.S., Hahn, E.M., Hermanns, H., Jansen, D.N.:
\newblock The ins and outs of the probabilistic model checker {MRMC}.
\newblock Performance Evaluation \textbf{68}(2) (2011)  90--104

\bibitem{prism_website}
{\tool{PRISM} Website}:
\newblock (2015) \url{http://prismmodelchecker.org}.

\bibitem{KNP12b}
Kwiatkowska, M., Norman, G., Parker, D.:
\newblock The {PRISM} benchmark suite.
\newblock  Proc.\ of QEST, IEEE CS (2012)  203--204

\bibitem{jani}
Budde, C.E., Dehnert, C., Hahn, E.M., Hartmanns, A., Junges, S., Turrini, A.:
\newblock Jani: Quantitative model and tool interaction.
\newblock (in preparation)

\bibitem{DBLP:conf/apn/EisentrautHK013}
Eisentraut, C., Hermanns, H., Katoen, J.P., Zhang, L.:
\newblock A semantics for every {GSPN}.
\newblock  Proc.\ of Petri Nets. Vol. 7927 of LNCS, Springer (2013)  90--109

\bibitem{handbook2002}
Stamatelatos, M., Vesely, W., Dugan, J.B., Fragola, J., Minarick, J.,
  Railsback, J.:
\newblock Fault Tree Handbook with Aerospace Applications.
\newblock NASA Headquarters (2002)

\bibitem{DBLP:series/mcs/McIverM05}
McIver, A., Morgan, C.:
\newblock Abstraction, Refinement and Proof for Probabilistic Systems.
\newblock Monographs in Computer Science. Springer (2005)

\bibitem{HJ94}
Hansson, H., Jonsson, B.:
\newblock A logic for reasoning about time and reliability.
\newblock Formal Aspects of Computing \textbf{6}(5) (1994)  512--535

\bibitem{DBLP:conf/cav/AzizSSB96}
Aziz, A., Sanwal, K., Singhal, V., Brayton, R.K.:
\newblock Verifying continuous time {M}arkov chains.
\newblock  Proc.\ of CAV. Vol. 1102 of LNCS, Springer (1996)  269--276

\bibitem{baier_cond}
Baier, C., Klein, J., Kl{\"{u}}ppelholz, S., M{\"{a}}rcker, S.:
\newblock Computing conditional probabilities in {M}arkovian models
  efficiently.
\newblock  Proc.\ of TACAS. Vol. 8413 of LNCS, Springer (2014)  515--530

\bibitem{DBLP:conf/atva/BrazdilCCFKKPU14}
Br{\'{a}}zdil, T., Chatterjee, K., Chmelik, M., Forejt, V.,
  Kret{\'{\i}}nsk{\'{y}}, J., Kwiatkowska, M., Parker, D., Ujma, M.:
\newblock Verification of {M}arkov decision processes using learning
  algorithms.
\newblock  Proc.\ of ATVA. Vol. 8837 of LNCS, Springer (2014)  98--114

\bibitem{eigenweb}
Guennebaud, G., Jacob, B.,  et~al.:
\newblock Eigen v3.
\newblock (http://eigen.tuxfamily.org)

\bibitem{gmmpp_website}
{\tool{gmm++} Website}:
\newblock (2015)
  \url{http://download.gna.org/getfem/html/homepage/gmm/index.html}.

\bibitem{cudd_website}
{\tool{CUDD} Website}:
\newblock (2015) \url{http://vlsi.colorado.edu/~fabio/CUDD/html/index.html}.

\bibitem{DBLP:conf/tacas/DijkP15}
van Dijk, T., van~de Pol, J.:
\newblock Sylvan: Multi-core decision diagrams.
\newblock  Proc.\ of TACAS. Vol. 9035 of LNCS, Springer (2015)  677--691

\bibitem{gurobi}
{Gurobi Optimization, Inc.}:
\newblock Gurobi optimizer reference manual.
\newblock http://www.gurobi.com (2015)

\bibitem{glpkweb}
{GNU project}:
\newblock Linear programming kit, version 4.6.
\newblock http://www.gnu.org/software/glpk/glpk.html (2016)

\bibitem{dMB08}
de~Moura, L.M., Bj{\o}rner, N.:
\newblock {Z3}: An efficient {SMT} solver.
\newblock  Proc.\ of TACAS. Vol. 4963 of LNCS, Springer (2008)  337--340

\bibitem{DBLP:conf/tacas/CimattiGSS13}
Cimatti, A., Griggio, A., Schaafsma, B.J., Sebastiani, R.:
\newblock The mathsat5 {SMT} solver.
\newblock  Proc.\ of TACAS. Vol. 7795 of Lecture Notes in Computer Science,
  Springer (2013)  93--107

\bibitem{Barrett10c.:the}
Barrett, C., Stump, A., Tinelli, C., Boehme, S., Cok, D., Deharbe, D.,
  Dutertre, B., Fontaine, P., Ganesh, V., Griggio, A., Grundy, J., Jackson, P.,
  Oliveras, A., Krstić, S., Moskal, M., Moura, L.D., Sebastiani, R., Cok,
  T.D., Hoenicke, J.:
\newblock C.: The smt-lib standard: Version 2.0.
\newblock Technical report (2010)

\bibitem{DBLP:journals/corr/abs-1305-5055}
Wimmer, R., Jansen, N., Vorpahl, A., {\'{A}}brah{\'{a}}m, E., Katoen, J.P.,
  Becker, B.:
\newblock High-level counterexamples for probabilistic automata.
\newblock Logical Methods in Computer Science \textbf{11}(1) (2015)

\bibitem{DBLP:journals/corr/DragerFK0U15}
Dr{\"{a}}ger, K., Forejt, V., Kwiatkowska, M., Parker, D., Ujma, M.:
\newblock Permissive controller synthesis for probabilistic systems.
\newblock Logical Methods in Computer Science \textbf{11}(2) (2015)

\bibitem{carl_website}
{\tool{CArL} Website}:
\newblock (2015) \url{http://goo.gl/8QsVxv}.

\bibitem{DBLP:conf/tacas/CalinescuJP16}
Calinescu, R., Johnson, K., Paterson, C.:
\newblock {FACT:} {A} probabilistic model checker for formal verification with
  confidence intervals.
\newblock  Proc.\ of TACAS. Vol. 9636 of LNCS, Springer (2016)  540--546

\bibitem{FIGpaper}
Budde, C.E., D'Argenio, P.R., Monti, R.E.:
\newblock Compositional construction of importance functions in fully automated
  importance splitting.
\newblock  Proc.\ of VALUETOOLS, {ICST} (2016)

\bibitem{DBLP:conf/atva/GuckTHRS14}
Guck, D., Timmer, M., Hatefi, H., Ruijters, E., Stoelinga, M.:
\newblock Modelling and analysis of {M}arkov reward automata.
\newblock  Proc.\ of ATVA. Vol. 8837 of LNCS, Springer (2014)  168--184

\bibitem{iscasmc}
Hahn, E.M., Li, Y., Schewe, S., Turrini, A., Zhang, L.:
\newblock {iscasMc}: {A} web-based probabilistic model checker.
\newblock  Proc.\ of FM. Vol. 8442 of LNCS, Springer (2014)  312--317

\bibitem{DBLP:conf/tacas/HartmannsH14}
Hartmanns, A., Hermanns, H.:
\newblock The modest toolset: An integrated environment for quantitative
  modelling and verification.
\newblock  Proc.\ of TACAS. Vol. 8413 of LNCS, Springer (2014)  593--598

\bibitem{PARAM10}
Hahn, E.M., Hermanns, H., Wachter, B., Zhang, L.:
\newblock {PARAM}: A model checker for parametric {M}arkov models.
\newblock  Proc.\ of CAV. Vol. 6174 of LNCS, Springer (2010)  660--664

\bibitem{DBLP:conf/tacas/HahnHWZ10}
Hahn, E.M., Hermanns, H., Wachter, B., Zhang, L.:
\newblock {PASS:} abstraction refinement for infinite probabilistic models.
\newblock  Proc.\ of TACAS. Vol. 6015 of LNCS, Springer (2010)  353--357

\bibitem{DBLP:conf/issre/LiuSD11}
Liu, Y., Sun, J., Dong, J.S.:
\newblock {PAT} 3: An extensible architecture for building multi-domain model
  checkers.
\newblock  Proc.\ of ISSRE, IEEE CS (2011)  190--199

\bibitem{KNP11}
Kwiatkowska, M., Norman, G., Parker, D.:
\newblock \textsc{Prism} 4.0: Verification of probabilistic real-time systems.
\newblock  Proc.\ of CAV. Vol. 6806 of LNCS, Springer (2011)  585--591

\bibitem{DBLP:journals/corr/abs-1207-1272}
Bulychev, P.E., David, A., Larsen, K.G., Mikucionis, M., Poulsen, D.B., Legay,
  A., Wang, Z.:
\newblock {UPPAAL-SMC:} statistical model checking for priced timed automata.
\newblock  Proc.\ of QAPL. Vol.~85 of {EPTCS} (2012)  1--16

\bibitem{jansen-et-al-atva-2012}
Jansen, N., {\'A}brah{\'a}m, E., Volk, M., Wimmer, R., Katoen, J.P., Becker,
  B.:
\newblock The {COMICS} tool -- {Computing} minimal counterexamples for {DTMCs}.
\newblock  Proc.\ of ATVA. Vol. 7561 of LNCS, Springer (2012)  349--353 (to
  appear).

\bibitem{ABBGS13}
Arnold, F., Belinfante, A., van~der Berg, F., Guck, D., Stoelinga, M.:
\newblock {DFTCalc}: {A} tool for efficient fault tree analysis.
\newblock  Proc.\ of SAFECOMP. Vol. 8153 of LNCS, Springer (2013)  293--301

\bibitem{DBLP:conf/apn/AmparoreBD14}
Amparore, E.G., Beccuti, M., Donatelli, S.:
\newblock (stochastic) model checking in {GreatSPN}.
\newblock  Proc.\ of Petri Nets. Vol. 8489 of LNCS, Springer (2014)  354--363

\end{thebibliography}

\documentclass{llncs}
\usepackage{amsmath,amssymb,amsfonts}
\usepackage{mathtools}
\usepackage[usenames,dvipsnames]{color}
\usepackage{xcolor}
\usepackage{xspace}
\usepackage{microtype}
\usepackage{todonotes}
\usepackage{colonequals}
\usepackage{marvosym}
\usepackage{wasysym}
\usepackage{booktabs}
\usepackage{refcount}
\usepackage{makecell}

\usepackage{multirow}
\usepackage{multicol}

\usepackage[T1]{fontenc}
\usepackage{paralist}

\usepackage{subfigure}
\usepackage{url}
\usepackage{appendix}
\usepackage{graphicx}
\usepackage{float}
\usepackage{algorithm}
\usepackage{listings}
\usepackage{newalg}
\usepackage{nicefrac}
\usepackage{tikz} 
\usepackage{pgfplots}
\usepackage{marvosym}
\usepackage{wrapfig}

\usepackage{footnote}
\usepackage{tablefootnote}
\usepackage{array}

\usetikzlibrary{arrows,decorations.pathmorphing,positioning,fit,trees,shapes,shadows,automata,calc} 

\usepackage{pdfpages}

\usepackage{preamble/macros} 

\begin{document}
  \title{The Probabilistic Model Checker \storm}
\subtitle{(Extended Abstract)}
\author{Christian Dehnert \and Sebastian Junges \and Joost-Pieter Katoen \and Matthias Volk}

\institute{RWTH Aachen University}
\vspace{1.4cm}
\maketitle

  \noindent We present a new probabilistic model checker \storm.
Using state-of-the-art libraries, we aim for both high performance and versatility.
This extended abstract gives a brief overview of the features of \storm.
  \paragraph{Goals.}
Probabilistic model checking has matured immensely in the past decade \cite{DBLP:conf/sfm/KwiatkowskaNP07,DBLP:conf/lics/Kwiatkowska03,DBLP:series/natosec/Baier16,Kat16}.
Applications and uses go far beyond the standard algorithms and objectives.
We therefore developed an extensible toolbox for probabilistic model checking that offers reusable modules to quickly implement new functionality.
The rise of more and more specialized tools witnesses that a single tool is rarely suitable for the broad range of applications and models.
Consequently, we offer a range of engines that make use of various backend solvers.
This core is aiming for high performance and low memory requirements.
On top of it we realized sophisticated techniques ranging from generation of probabilistic counterexamples \cite{DBLP:conf/atva/DehnertJWAK14}, over permissive scheduler generation \cite{junges-et-al-tacas-2016} and parameter synthesis \cite{dehnert-et-al-cav-2015,DBLP:conf/atva/QuatmannD0JK16} to the analysis of dynamic fault trees \cite{DBLP:conf/safecomp/0001JK16} and probabilistic program verification \cite{DBLP:conf/atva/0001DKKW16}.
Finally, the wide range of input languages supports the development of new techniques in a variety of fields.
\paragraph{Supported model types.} \storm{} supports both discrete-time and con\-tin\-u\-ous-time Markov models and nondeterministic variants thereof. 
Table~\ref{tab:modeltypes} gives an overview and classification about the concrete model types: discrete-time Markov chains (DTMCs), continuous-time Markov chains (CTMCs), Markov decision processes (MDPs) and Markov automata (MA), the compositional variant of continuous-time Markov Decision processes (CTMDPs) \cite{Kat16}.
Additionally, all these model types can be enriched with reward models.
\begin{table}
  \def\arraystretch{1.3}\tabcolsep=10pt
  \begin{center}
    \caption{Model types supported by \storm.}
    \begin{tabular}{ccc}
      \toprule
      & discrete-time & continuous-time \\
      \cmidrule{1-3}
      deterministic & DTMCs & CTMCs \\
      nondeterministic & MDPs & MA \\
      \bottomrule
    \end{tabular} 
    \label{tab:modeltypes}
  \end{center}
\end{table}
\paragraph{Modelling languages.}
There are several modelling languages that can be used to specify the aforementioned model types. 
First of all, in the spirit of MRMC \cite{mrmc}, models can be passed in a format that explicitly enumerates transitions. 
However, \storm{} also supports a variety of symbolic input formats. 
Most prominently, the \prism{} input language \cite{prism_website} is supported, which enables us to consider all benchmark models from the \prism{} benchmark suite \cite{KNP12b} that represent supported model types. 
In an attempt to unify the probabilistic modelling language landscape, the JANI format \cite{jani} is currently under development and its first version is included in \storm.
Besides, it is the first tool to support \emph{every} generalized stochastic Petri net (GSPN) \cite{DBLP:conf/apn/EisentrautHK013} via both a dedicated model builder as well as an encoding in JANI.
Additionally, \storm{} features the analysis of dynamic fault trees (DFTs) \cite{handbook2002} and has been shown to outperform competing tools in this domain \cite{DBLP:conf/safecomp/0001JK16}.
Finally, we support probabilistic programs written in the conditional probabilistic guarded command language (cpGCL) \cite{DBLP:series/mcs/McIverM05}, again via a reduction to JANI.

\paragraph{Properties.}
The main focus of our tool is probabilistic branching time logics, i.e.~PCTL \cite{HJ94} and CSL \cite{DBLP:conf/cav/AzizSSB96} for discrete-time and continuous-time models, respectively. 
To enable the treatment of reward-objectives, we support reward extensions of these logics in a similar way as PRISM.
As conditional probabilities and conditional rewards \cite{baier_cond} have proven to express interesting properties, we also support these.
%
\paragraph{Engines.}
There are different representations of probabilistic models in memory that differ in efficiency depending on the characteristics of the input model.
\storm{} features two distinct representations: while \emph{sparse matrices} tend to work well for small and moderately sized models, multi-terminal decision diagrams (MTBDDs) are able to represent gigantic systems.
To enable the treatment of a broader class of input models, the user can select between several engines built around the two in-memory representations to perform the verification tasks.
Both \storm's \textit{sparse} and the \textit{learning} engine purely use a sparse matrix based representation. While the former amounts to an efficient implementation of the standard approaches, the latter one implements the ideas of \cite{DBLP:conf/atva/BrazdilCCFKKPU14} to provide sound statistical model checking for discrete-time models.
Two other engines, \textit{dd} and \textit{hybrid}, use MTBDDs as the primary representation.
While \textit{dd} exclusively uses decision diagrams, \textit{hybrid} also uses sparse matrices for operations deemed more suitable on this data format.
\paragraph{Solvers.}
\storm's infrastructure is built around the notion of a \emph{solver}. 
For instance, solvers are available for sets of linear or Bellman equations (both using sparse matrices as well as MTBDDs), (mixed-integer) linear programming (MILP) and satisfiability modulo theories (SMT) solving. 
Offering these interfaces has several key advantages.
First, it provides easy and coherent access to the tasks commonly involved in probabilistic model checking.
Secondly, it enables the use of dedicated state-of-the-art high-performance libraries for the task at hand. More specifically, as the performance characteristics of different backend solvers can vary drastically for the same input, this permits choosing the best solver for a given task.
Licensing problems are avoided, because implementations can be easily enabled and disabled, depending on whether or not the particular license fits the requirements.
Finally, implementing new solver functionality is easy and can be done without knowledge about the global code base.
For each of those interfaces, several actual implementations exist. Table~\ref{tab:solvers} gives an overview over the currently available implementations.
\begin{table}[t]
  \def\arraystretch{1.3}\tabcolsep=10pt
  \begin{center}
  
    \caption{Solvers offered by \storm.}
    \begin{tabular}{ll}
      \toprule
      solver type & available solvers \\
      \cmidrule{1-2}
      linear equations (sparse) & Eigen \cite{eigenweb}, gmm++ \cite{gmmpp_website}, built-in \\
      linear equations (MTBDD) & CUDD \cite{cudd_website}, Sylvan \cite{DBLP:conf/tacas/DijkP15} \\
      Bellman equations (sparse) & Eigen, gmm++, built-in \\
      Bellman equations (MTBDD) & CUDD, Sylvan \\
      (MI)LP & Gurobi \cite{gurobi}, glpk \cite{glpkweb} \\
      SMT & Z3 \cite{dMB08}, MathSAT \cite{DBLP:conf/tacas/CimattiGSS13}, SMTLIB \cite{Barrett10c.:the} \\
      \bottomrule
    \end{tabular} 
    \label{tab:solvers}
  \end{center}
\end{table}
Many components in \storm{} make use of these solvers. 
The most prominent example is the obvious use of the equation solvers for answering standard verification queries. 
However, various other modules use them too, for example model generation (SMT), counterexample generation \cite{DBLP:journals/corr/abs-1305-5055,DBLP:conf/atva/DehnertJWAK14} (SMT, MILP) and permissive scheduler generation \cite{DBLP:journals/corr/DragerFK0U15,junges-et-al-tacas-2016} (SMT, MILP).

\paragraph{Parametric models and exact arithmetic.}
\storm was used as backend in \cite{dehnert-et-al-cav-2015,DBLP:conf/atva/QuatmannD0JK16}. By using the dedicated library \tool{CArL} \cite{carl_website} for the representation of rational functions and applying novel algorithms for the analysis of parametric discrete-time models, it has proven to significantly outperform other tools. 
In addition, for these models \storm{} is able to compute exact solutions rather than using floating point arithmetic.

\paragraph{Usage.}
\storm{} can be used via three interfaces. 
For end-users we provide command-line interfaces: There are several binaries that provide specialized access to the available settings for different tasks. 
For example, \tool{storm-dft} offers the features and settings related to the analysis of DFTs. 
Advanced users can utilize one of the many settings to tune the performance.
Developers may either use a C++ API that offers fine-grained and performance-oriented access to \storm's functionality, or use a Python API which supports rapid prototyping and allows users to profit from high-performance implementations within \storm.

\paragraph{Related work.}
There are several tools whose functionality overlaps with that of \storm (in alphabetical order): \tool{FACT} \cite{DBLP:conf/tacas/CalinescuJP16}, \tool{FIG} \cite{FIGpaper}, \tool{IMCA} \cite{DBLP:conf/atva/GuckTHRS14}, \tool{iscasMC} \cite{iscasmc}, \tool{Modest} \cite{DBLP:conf/tacas/HartmannsH14}, \tool{MRMC} \cite{mrmc}, \tool{PARAM} \cite{PARAM10}, \tool{PASS} \cite{DBLP:conf/tacas/HahnHWZ10}, \tool{PAT} \cite{DBLP:conf/issre/LiuSD11}, \tool{PRISM} \cite{KNP11} and \tool{Uppaal SMC} \cite{DBLP:journals/corr/abs-1207-1272}. Moreover, there are specialized tools for the generation of probabilistic counterexamples (\tool{COMICS} \cite{jansen-et-al-atva-2012}), DFTs (\tool{DFTCalc} \cite{ABBGS13}) and GSPNs (\tool{GreatSPN} \cite{DBLP:conf/apn/AmparoreBD14}).
Despite the existence of these tools, we believe that \storm{} is unique in its trade-off between performance and modularity, the supported solvers and wide range of supported modelling languages.

  \pagebreak
  \bibliographystyle{splncs}
  \bibliography{literature}
\end{document}